OpenStudy (anonymous):

Which of the following shows the right side of the equation after applying a reasonable first step to solve negative h divided by two plus negative one-fourth equals four-ninths? four-ninths minus one-fourth four-ninths plus four-ninths four-ninths plus one-fourth four-ninths minus four-ninths

OpenStudy (anonymous):

@texaschic101

OpenStudy (texaschic101):

-h/2 + (-1/4) = 4/9 -h/2 - 1/4 = 4/9 -h/2 = 4/9 + 1/4 <=== answer 3